# SyncTwin: Fast Digital Twin Construction and Synchronization for Safe Robotic Grasping

### v1

@@ Abstract
Accurate and safe grasping under dynamic and visually occluded conditions remains a core challenge in real-world robotic manipulation. We introduce SyncTwin, a novel digital twin framework that unifies fast 3D scene reconstruction and real-to-sim synchronization for robust and safety-aware grasping in such environments. In the offline stage, we employ VGGT to rapidly reconstruct object-level 3D assets from RGB images, forming a reusable geometry library for simulation. During execution, SyncTwin continuously synchronizes the digital twin by tracking real-world object states via point cloud segmentation updates and aligning them through colored-ICP registration. The updated twin enables motion planners to compute collision-free and dynamically feasible trajectories in simulation, which are safely executed on the real robot through a closed real-to-sim-to-real loop. Experiments in dynamic and occluded scenes show that SyncTwin improves grasp accuracy and motion safety, demonstrating the effectiveness of digital-twin synchronization for real-world robotic execution.
1 Introduction
@@
Achieving accurate and safe robotic grasping in dynamic real-world environments remains a long-standing challenge, due to incomplete perception and dynamic scenes. Without an accurate understanding of their surroundings, robots risk colliding with the environment, which may damage hardware or even endanger humans [58]. Thus, ensuring that robots can safely plan and execute motions amid dynamic scene changes is a prerequisite for reliable real-world deployment. Unlike simulation, where complete scene geometry and object states are fully accessible, real-world perception often only offers partial, occluded observations. As a result, motion-planning algorithms [33, 16, 36] that rely on full and accurate environment models in simulation face significant challenges when deployed in the real world.
@@
Several recent efforts have attempted to bridge the gap between perception and control. Voxel-based mapping systems like NVBlox [23] enable online obstacle reconstruction, yet they only provide a voxel grid, which is a representation too coarse for reliable manipulation. End-to-end reactive policies like DRP [53] handle dynamics through continuous control but require robot-specific retraining and lack generalization across new hardware or environments. Thus, all of these methods share a critical limitation: they operate without a consistent and complete model of the scene, leading to unsafe or unreliable executions in the real world.
@@
Bridging this gap between simulation and the real world, a robot needs a perception model that can efficiently perceive what objects exist in the environment and also track how the real-world scene evolves over time. In other words, instead of planning in a static or outdated simulation, the robot should plan within a dynamic digital twin that mirrors the physical world in real time, where perception and control are tightly coupled through continuous real-to-sim synchronization. However, real-world perception is inherently partial—single-view occlusions often reveal only fragments of object geometry, making grasping and motion planning unreliable. Inspired by SAM4D [51], which maintains a memory bank of object assets, we incorporate the idea of leveraging object-level memories to complete partial observations at execution time. We develop SyncTwin, a digital-twin framework equipped with an object memory bank that performs real-time object tracking from point clouds, injects accurate poses and geometries into simulation, and closes the loop by executing planned trajectories back on the real robot, as illustrated in Figure LABEL:fig:overview.
@@
SyncTwin operates in two stages. In Stage I, VGGT [46] generates scene-level point clouds from RGB inputs. Object-level point clouds are extracted via projection, segmentation, and denoising, then converted into lightweight meshes and stored in a memory bank. In Stage II, the digital twin is continuously synchronizes with the real world using SAM2 for object tracking and GPU-accelerated colored-ICP [42]. Combined, these two methods align partial observations with stored assets to maintain a consistent and complete scene representation. The updated digital twin is streamed into Isaac Sim [24], allowing cuRobo [41] to perform motion planning and generate collision-aware trajectories. Moreover, this architecture can be transferred to different real robots without any retraining.
@@
Figure 2: Framework of the SyncTwin. Stage I reconstructs simulation-ready 3D assets from RGB images using VGGT and SAM2. Multi-view masks are unprojected into point clouds, then denoised, scaled, and meshed into clean object assets stored in a memory bank. Stage II performs real-time object segmentation, pose tracking, and asset-based completion, enabling grasp generation and reactive motion planning in a closed real2sim2real loop. By continuously updating the digital twin and leveraging simulation for decision making, the system ensures safe and adaptive execution under dynamic and partially occluded environments.
2 Related Work
3D Scene Reconstruction and Segmentation. Traditional 3D scene reconstruction approaches rely on RGB-D input from depth sensors to estimate geometry [28, 38, 7, 5]. More recent approaches operate purely on RGB images, either through optimization-based pipelines such as Structure-from-Motion (SfM) [30, 35, 31, 14] and Multi-View Stereo (MVS) [37, 10, 11], or through learning-based frameworks such as MVSNet [54], NeRF [22], and 3D Gaussian Splatting [20]. While these methods can recover visually plausible geometry, they often require substantial computation time to produce complete reconstructions. For robotics applications, object-centric 3D segmentation is important as well. Prior work has explored segmentation directly on point clouds using learned clustering and aggregation strategies [57, 47, 15], as well as multi-view projection of 2D masks for 3D instance segmentation [34, 21, 56]. However, projection-based methods typically require accurate camera extrinsics [52, 2], and little research has examined how such strategies perform when applied to point clouds produced by learning-based reconstruction methods. Our work addresses this gap by unifying learned 3D reconstruction with mask-guided multi-view segmentation, enabling efficient and robust object asset generation for downstream robotic manipulation.
@@ 2 Related Work
Safe Motion Planning for Manipulation. Safe and robust motion planning remains a critical challenge for robotic manipulation in unstructured environments [39, 18]. Classical and learning-based planners [33, 16, 36] typically operate in simulation, where the environment is fully accessible. When deployed in the real world, however, they must handle partial observations. To narrow this gap, several methods construct static point cloud maps and import them into simulation for offline planning [3], though such maps struggle to support real-time adaptation. Other approaches aim to ensure safety by predicting collision-free actions directly from images in latent space [27, 1], using force-sensing-based control [48, 17], or distilling planning policies from point cloud observations [6, 53, 9]. NVBlox improves online safety by voxelizing scene geometry and segmenting the robot in real time [23]. In contrast, our method maintains a dynamically synchronized digital twin that continuously provides updated scene geometry to the planner, enabling safe execution in dynamic, cluttered environments.
3 Method
@@
SyncTwin consists of two stages: (1) fast digital twin construction, and (2) digital twin synchronization. An overall framework is provided in Figure 2.
3.1 Problem Formulation
@@
We aim to enable safe robotic grasping in dynamic, partially observable real-world environments by maintaining a continuously synchronized digital twin. This problem can be decomposed into the following components:
@@
Stage I: Fast Digital Twin Construction. The system receives a small set of RGB images $\{\mathbf{I}_{i}\}_{i=1}^{N}$ along with camera intrinsics $K$ and estimated extrinsics $\{\mathbf{T}_{i}^{\text{world}}\}_{i=1}^{N}$. The goal is to produce object-level, simulation-ready 3D assets $\mathbb{B}=\{\mathcal{X}_{j},\mathcal{M}_{j}^{3D}\}$ from these images. The main challenge is that learning-based extrinsics contain unstable errors, causing mask–projection misalignment and table–object mixing in the reconstructed point cloud, which must be addressed to obtain clean object geometry.
@@
Stage II: Online Digital Twin Synchronization. During execution, the system receives streaming RGB-D frames and corresponding partial point clouds $\mathcal{X}_{p}$. The objective is to maintain accurate object poses $\mathbf{T}_{j}^{\text{world}}$ in the simulator by aligning $\mathcal{X}_{p}$ with their complete assets $\mathcal{X}_{m}\in\mathbb{B}$. This synchronized scene is streamed into Isaac Sim [24], where cuRobo’s MPC planner [41] produces short-horizon, collision-free trajectories $\mathbf{A}_{t:t+H}=\{\mathbf{a}_{0},\ldots,\mathbf{a}_{H}\}$. The key challenge is robustly tracking objects under occlusion and partial observation, while ensuring that the object poses and geometries can be accurately updated into the simulator to enable safe real-to-sim-to-real planning.
3.2 Fast Digital Twin Construction
@@
The first stage aims to rapidly reconstruct the 3D environment and extract object-level representations suitable for real-time simulation, where the focus is to achieve accurate geometric perception for motion planning, rather than photorealistic reconstruction. We employ VGGT [46] to reconstruct dense scene point clouds directly from a small number of RGB images, which enables efficient extraction of object-level geometry in a fast and low-cost manner without depth sensors or multi-view optimization.
@@
Nevertheless, there are two major practical challenges. First, the camera extrinsics estimated by VGGT are often inaccurate, which leads to projection misalignment during mask-based segmentation. Second, the generated point cloud is not in a true metric world scale, causing the imported assets to appear incorrectly sized relative to the robot in the simulator. Therefore, Stage I focuses on producing simulation-ready 3D digital assets from VGGT outputs by addressing these limitations. We design a four-step pipeline that corrects extrinsic inaccuracies, enforces scale consistency, and generates clean object meshes suitable for downstream planning, which is introduced as follows.
@@
Mask Projection Expansion. To mitigate inaccuracies in VGGT-estimated camera extrinsics, each 2D segmentation mask $\mathcal{S}_{i}$ is spatially expanded before projection, which ensures full coverage of object boundaries and prevents missing edge regions during 3D reconstruction. While mask expansion compensates for projection drift, it also introduces floating outliers (e.g., background points above the object) and merged support-plane regions (e.g., table surfaces). To address this, we apply our point clouds denoising mechanism to isolate the true object shape.
@@
Figure 3: Supporting-plane noise removal mechanism. A virtual light sphere expands from the object center to identify openings and boundary points, enabling filtering of table-plane noise.
@@
Point Clouds Denoising. Detecting openings or cavities in 3D point clouds is a fundamental step in shape understanding for denoising table noise. We propose a purely geometric method that progressively expands a virtual light sphere from the object’s center and tracks uncovered regions on the spherical sampling space. The algorithm automatically detects openings and extracts rim points around their boundaries without requiring mesh topology or prior segmentation. Figure 3 shows the overall process.
@@
Given a point cloud $\mathcal{P}$, we first estimate a geometric center $\mathbf{c}=\mathrm{mean}(\mathcal{P})$. We then discretize the unit sphere into $F$ directions $\{\mathbf{d}_{i}\}_{i=1}^{F}$ using a Fibonacci spiral distribution [13], forming a uniform sampling domain $\mathcal{D}$. Each point $\mathbf{p}_{i}$ defines a normalized direction $\hat{\mathbf{v}}_{i}=(\mathbf{p}_{i}-\mathbf{c})/\|\mathbf{p}_{i}-\mathbf{c}\|$ and is assigned to its nearest angular bucket direction $\mathbf{d}_{j}$ if $\hat{\mathbf{v}}_{i}\cdot\mathbf{d}_{j}\geq\cos(\theta_{\text{tolerance}})$, We then iteratively expand a virtual sphere centered at $\mathbf{c}$ with radius $r_{t}$:
@@
A direction $\mathbf{d}_{j}$ is marked as hit once any assigned point enters the sphere, and unhit directions form a binary mask $\mathcal{U}_{t}$. Stable uncovered regions (openings or cavities) are detected when the largest unhit component remains consistent over iterations. We denote by $\mathcal{N}(j)=\{k\mid(j,k)\in\mathcal{P}\}$ the neighborhood of bucket $j$. Boundary buckets are defined as unhit directions adjacent to hit ones:
@@
For each boundary $\mathbf{d}_{j}$, the farthest point within tolerance is chosen as a rim sample $\mathbf{p}^{*}_{j}$. Connected components on the spherical adjacency graph are extracted to identify large uncovered regions. Let $\{\mathcal{C}_{m}\}$ denote all connected components of the unhit set $\mathcal{U}_{t},\mathcal{C}_{\max}$ is the largest connected region. The principal opening direction is then computed by averaging the largest unhit component: $\mathbf{n}_{\text{open}}=\frac{\sum_{j\in\mathcal{C}_{\max}}\mathbf{d}_{j}}{\|\sum_{j\in\mathcal{C}_{\max}}\mathbf{d}_{j}\|}.$ Finally, the rim points $\{\mathbf{p}^{*}_{j}\}$ are fitted with a plane using SVD [12], yielding the opening orientation and visualizable boundary. Compared with RANSAC plane fitting [8], which can only segment dominant planes, our method can detect cavity openings, enabling the identification of ring-shaped planes surrounding the opening
@@
Real-World Scale Alignment. Since VGGT produces point clouds that are not in a world-scale metric, we estimate a global scale factor to align the reconstructions with world coordinates. Even with known intrinsic parameters, according to the pinhole camera model, monocular geometry cannot determine absolute scale [14]. Therefore, we calibrate the scale using a reference object or markers of known physical dimensions within the scene. Implementation details are provided in supplementary.
@@
Mesh Simplification. To maintain real-time performance in the digital twin, we apply an adaptive mesh decimation that reduces vertex count while preserving geometric fidelity and collision boundaries. To avoid over-smoothing across sharp edges, we use an angle-based gating weight $w_{ij}=1$ if $\theta_{ij}\leq\theta_{\mathrm{th}}$ and 0 otherwise, where $\theta_{\mathrm{th}}$ is a feature threshold (e.g., $30^{\circ}$) and $\theta_{ij}=\arccos(\mathbf{n}_{i}^{\top}\mathbf{n}_{j})$ for all $j\in\mathcal{N}(i)$. Each vertex is then updated via a selective Laplacian step [44]. Compared to uniform mesh decimation, this feature-aware smoothing preserves sharp edges around handles and object rims, which are critical for accurate grasp planning. All processed point clouds, meshes, and their id are stored in a memory bank, which serves for object recognition and scene synchronization in Stage II.
3.3 Online Digital Twin Synchronization

### v2

# SyncTwin: Fast Digital Twin Construction and Synchronization for Safe Robotic Manipulation

@@ Abstract
Accurate and safe robotic manipulation under dynamic and visually occluded conditions remains a core challenge in real-world deployment. We introduce SyncTwin, a novel digital twin framework that unifies fast 3D scene reconstruction and real-to-sim synchronization for robust and safety-aware robotic manipulation in such environments. In the offline stage, we employ VGGT to rapidly reconstruct object-level 3D assets from RGB images, forming a reusable geometry library. During execution, SyncTwin continuously synchronizes the digital twin by tracking real-world object states via point cloud segmentation updates and aligning them through colored-ICP registration. The synchronized twin enables motion planners to compute collision-free and dynamically feasible trajectories in simulation, which are safely executed on the real robot through a closed real-to-sim-to-real loop. Experiments in dynamic and occluded scenes show that SyncTwin improves manipulation performance and motion safety, demonstrating the effectiveness of digital twin synchronization for real-world robotic execution. The video demos and code can be found on the project website: https://sync-twin.github.io/.
I Introduction
@@
Achieving accurate and safe robotic manipulation in dynamic real-world environments remains a long-standing challenge due to incomplete perception and scene dynamics. Without an accurate understanding of their surroundings, robots risk colliding with the environment, which may damage hardware or even endanger humans [1]. Thus, both in real-robot online training and in real-world robot model deployment, enabling robots to safely plan and execute motions under dynamic scene changes is a critical concern. Unlike simulation, where complete scene geometry and object states are fully accessible, real-world perception often provides only partial and occluded observations. As a result, although many motion-planning algorithms [2, 3, 4] demonstrate strong performance in simulation, robust motion planning in real-world environments remains an open problem with significant challenges.
@@
Several recent efforts have attempted to bridge the gap between perception and control. Voxel-based mapping systems like NVBlox [5] enable online obstacle reconstruction, yet they only provide a voxel grid, which is a representation too coarse for reliable manipulation. End-to-end reactive policies like DRP [6] handle dynamics through continuous control but require robot-specific retraining and lack generalization across new hardware or environments. And both share a critical limitation: they operate without a consistent and complete model of the scene, leading to unsafe or unreliable executions.
@@
Bridging this gap between simulation and the real world, a robot needs a perception model that can efficiently perceive what objects exist in the environment and also track how the real-world scene evolves over time. In other words, instead of planning in a static or outdated simulation, the robot should plan within a dynamic digital twin that mirrors the physical world in real time, where perception and control are tightly coupled through continuous real-to-sim synchronization. However, real-world perception is inherently partial single-view occlusions often reveal only fragments of object geometry, making grasping and motion planning unreliable. Inspired by SAM4D [7], which maintains a memory bank of object assets, we incorporate the idea of leveraging object-level memories to complete partial observations at execution time. We develop SyncTwin, a digital twin framework equipped with an object memory bank that performs real-time object tracking from point clouds, injects accurate poses and geometries into simulation, and closes the loop by executing planned trajectories back on the real robot, addressing a key limitation of prior digital twin systems, which do not support continuous real-to-sim-to-real synchronization, as illustrated in Fig. LABEL:fig:overview.
@@
SyncTwin operates in two stages. In Stage I, VGGT [8] generates scene-level point clouds from RGB inputs. Object-level point clouds are extracted via projection, segmentation, and denoising, then converted into lightweight meshes and stored in a memory bank. In Stage II, the digital twin is continuously synchronized with the real world through online object tracking and GPU-accelerated colored-ICP [9]. Combined, these two methods align partial observations with stored assets to maintain a consistent and complete scene representation. The updated digital twin is streamed into Isaac Sim [10], allowing cuRobo [11] to perform motion planning and generate collision-aware trajectories. Moreover, this architecture can be transferred to different real robots without any retraining.
@@
Figure 2: Framework of the SyncTwin. Stage I constructs simulation-ready 3D assets from RGB images using VGGT. Multi-view masks are unprojected into point clouds, then denoised, scaled, and meshed into clean object assets stored in the memory bank. Stage II performs real-time object segmentation, pose tracking, and asset-based completion, enabling grasp generation and reactive motion planning in a closed real2sim2real loop. By continuously updating the digital twin and leveraging simulation for decision making, the system ensures safe and adaptive execution under dynamic and partially occluded environments.
II Related Work
3D Scene Reconstruction and Segmentation. Traditional 3D scene reconstruction approaches rely on RGB-D input from depth sensors to estimate geometry [12, 13, 14, 15]. More recent approaches operate purely on RGB images, either through optimization-based pipelines such as Structure-from-Motion (SfM) [16, 17, 18, 19] and Multi-View Stereo (MVS) [20, 21, 22], or through learning-based frameworks such as MVSNet [23], NeRF [24], and 3D Gaussian Splatting [25]. While these methods can recover visually plausible geometry, they often require substantial computation time to produce complete reconstructions. For robotics applications, object-centric 3D segmentation is important as well. Prior work has explored segmentation directly on point clouds using learned clustering and aggregation strategies [26, 27, 28], as well as multi-view projection of 2D masks for 3D instance segmentation [29, 30, 31]. However, projection-based methods typically require accurate camera extrinsics [32, 33], and little research has examined how such strategies perform when applied to point clouds produced by learning-based reconstruction methods. Our work addresses this gap by unifying learned 3D reconstruction with mask-guided multi-view segmentation to achieve efficient and robust object asset generation, providing geometry-aware perception to support safe and reliable motion planning in robotic manipulation.
@@ II Related Work
Safe Motion Planning for Manipulation. Safe and robust motion planning remains a critical challenge for robotic manipulation in unstructured environments [42, 43]. Classical and learning-based planners [2, 3, 4] typically operate in simulation, where the environment is fully accessible. When deployed in the real world, however, they must handle partial observations. To narrow this gap, several methods construct static point cloud maps and import them into simulation for offline planning [44], though such maps struggle to support real-time adaptation. Other approaches aim to ensure safety by predicting collision-free actions directly from images in latent space [45, 46], using force-sensing-based control [47, 48], or distilling planning policies from point cloud observations [49, 6, 50]. NVBlox improves online safety by voxelizing scene geometry and segmenting the robot in real time [5]. In contrast, our method maintains a dynamically synchronized digital twin that continuously provides updated scene geometry to the planner, enabling safe execution in dynamic environments.
III Method
@@
SyncTwin consists of two stages: (1) fast digital twin construction and (2) digital twin synchronization. An overall framework of the proposed method is provided in Fig 2.
III-A Problem Formulation
@@
We aim to enable safe robotic manipulation in dynamic, partially observable real-world environments by maintaining a continuously synchronized digital twin. This problem can be decomposed into the following components:
@@
Stage I: Fast Digital Twin Construction. Given a small set of RGB images $\{\mathbf{I}_{i}\}_{i=1}^{N}$ along with intrinsics $K$ and estimated extrinsics $\{\mathbf{T}_{i}^{\text{world}}\}_{i=1}^{N}$. Our goal is to generate object-level, simulation-ready 3D assets $\mathbb{B}=\{\mathcal{X}_{j},\mathcal{M}_{j}^{3D}\}$ from these images. The main challenge is that learning-based extrinsics contain unstable errors, causing mask-projection misalignment and table-object mixing in the reconstructed point cloud, which must be addressed to obtain clean object geometry.
@@
Stage II: Online Digital Twin Synchronization. During execution, the fixed calibrated camera receives streaming RGB-D frames and corresponding partial point clouds $\mathcal{X}_{p}$. The objective is to maintain accurate object poses $\mathbf{T}_{j}^{\text{world}}$ in the simulator by aligning $\mathcal{X}_{p}$ with their complete assets $\mathcal{X}_{m}\in\mathbb{B}$. This synchronized scene is streamed into Isaac Sim [10], where cuRobo’s MPC planner [11] produces short-horizon, collision-free trajectories $\mathbf{A}_{t:t+H}=\{\mathbf{a}_{0},\ldots,\mathbf{a}_{H}\}$. The key challenge is robustly tracking objects under occlusion and partial observation, while ensuring that the object poses and geometries can be accurately updated into the simulator to enable safe real-to-sim-to-real planning and execution.
III-B Fast Digital Twin Construction
@@
The first stage aims to rapidly reconstruct the 3D environment and extract object-level representations suitable for real-time simulation, where the focus is to achieve accurate geometric perception for motion planning, rather than photorealistic reconstruction. We employ VGGT [8] to reconstruct dense scene point clouds directly from a small number of RGB images, which enables efficient extraction of object-level geometry in a fast and low-cost manner without relying on multiple depth sensors or multi-view optimization, while avoiding repeated calibration of the Realsense’s extrinsic.
@@
Nevertheless, there are two major practical challenges. First, the camera extrinsics estimated by VGGT are often inaccurate, which leads to projection misalignment during mask-based segmentation. Second, the generated point cloud is not in a true metric world scale, causing the imported assets to appear incorrectly sized relative to the robot in the simulator. Therefore, Stage I focuses on producing simulation-ready 3D digital assets from VGGT outputs by addressing these limitations. We design a four-step pipeline that solves extrinsic inaccuracies, scale consistency, and generates clean object meshes suitable for planning, which is introduced as follows.
@@
Mask Projection Expansion. To mitigate inaccuracies in VGGT-estimated camera extrinsics, each 2D segmentation mask $\mathcal{S}_{i}$ is spatially expanded before projection, which ensures full coverage of object boundaries and prevents missing edge regions. While mask expansion compensates for projection drift, it also introduces floating outliers (e.g., background points around the object) and merged support-plane regions (e.g., table surfaces). To address this, we apply our point clouds denoising mechanism to isolate the true object shape.
@@
Figure 3: Supporting-plane noise removal mechanism. A virtual light sphere expands from the object center to identify openings and boundary points, enabling filtering of table noise.
@@
Point Clouds Denoising. Detecting openings or cavities in 3D point clouds is a fundamental step in shape understanding for denoising table noise. We propose a purely geometric method that progressively expands a virtual light sphere from the object’s center and tracks uncovered regions on the spherical sampling space. The algorithm automatically detects openings and extracts rim points around their boundaries without manual boundary selection. Fig. 3 shows the overall process.
@@
Given a point cloud $\mathcal{P}$, we first estimate a geometric center $\mathbf{c}=\mathrm{mean}(\mathcal{P})$. We then discretize the unit sphere into $F$ directions $\{\mathbf{d}_{i}\}_{i=1}^{F}$ using a Fibonacci spiral distribution [51], forming a uniform sampling domain $\mathcal{D}$. Each point $\mathbf{p}_{i}$ defines a normalized direction $\hat{\mathbf{v}}_{i}=(\mathbf{p}_{i}-\mathbf{c})/\|\mathbf{p}_{i}-\mathbf{c}\|$ and is assigned to its nearest angular bucket direction $\mathbf{d}_{j}$ if $\hat{\mathbf{v}}_{i}\cdot\mathbf{d}_{j}\geq\cos(\theta_{\text{tolerance}})$. We then iteratively expand a virtual sphere centered at $\mathbf{c}$ with radius $r_{t}$, along the sampled directions.
@@
A direction $\mathbf{d}_{j}$ is marked as hit once any assigned point enters the sphere, and unhit directions form a binary mask $\mathcal{U}_{t}$. The largest unhit connected component is identified as an opening or cavity after iterations. We denote by $\mathcal{N}(j)=\{k\mid(j,k)\in\mathcal{P}\}$ the neighborhood of bucket $j$. Boundary buckets are defined as unhit directions adjacent to hit ones:
@@
For each boundary $\mathbf{d}_{j}$, the farthest point within tolerance is chosen as a rim sample $\mathbf{p}^{*}_{j}$. Connected components on the spherical adjacency graph are extracted to identify large uncovered regions. Let $\{\mathcal{C}_{m}\}$ denote all connected components of the unhit set $\mathcal{U}_{t}$, $\mathcal{C}_{\max}$ is the largest connected region. The principal opening direction is then computed by averaging the largest unhit component: $\mathbf{n}_{\text{open}}=\frac{\sum_{j\in\mathcal{C}_{\max}}\mathbf{d}_{j}}{\|\sum_{j\in\mathcal{C}_{\max}}\mathbf{d}_{j}\|}.$ Finally, the rim points $\{\mathbf{p}^{*}_{j}\}$ are fitted with a plane using SVD [52] to remove the supporting surface noise points. Compared with RANSAC plane fitting [53], which can only segment dominant planes, our method can detect cavity openings, enabling the identification of ring-shaped planes surrounding the opening.
@@
Real-World Scale Alignment. Since VGGT produces point clouds that are not in a world-scale metric, we estimate a global scale factor to align the reconstructions with world coordinates. Even with known intrinsic parameters, according to the pinhole camera model, monocular geometry cannot determine absolute scale [19]. Therefore, we calibrate the scale using a marker of known physical dimensions within the scene. Implementation details are provided in the Appendix.
@@
Mesh Simplification. To maintain real-time performance in the digital twin, we apply an adaptive mesh decimation that reduces vertex count while preserving geometric fidelity. To avoid over-smoothing across sharp edges, we use an angle-based gating weight $w_{ij}=1$ if $\theta_{ij}\leq\theta_{\mathrm{th}}$ and 0 otherwise, where $\theta_{\mathrm{th}}$ is a feature threshold (e.g., $30^{\circ}$) and $\theta_{ij}=\arccos(\mathbf{n}_{i}^{\top}\mathbf{n}_{j})$ for all $j\in\mathcal{N}(i)$. Each vertex is then updated via a selective Laplacian step [54]. Compared to uniform mesh decimation, this feature-aware smoothing preserves sharp edges around handles and object rims, which are critical for accurate grasp planning. All processed point clouds, meshes, and their id are stored in a memory bank, which serves for scene synchronization in Stage II.
III-C Online Digital Twin Synchronization